OpenStudy (anonymous):

Vectors Question! If 2x+3y=a and -x+5y=6b, express x and y in terms of a and b.

OpenStudy (kainui):

Can you throw this into a matrix equation? Then once you've done that, just find the inverse matrix and that'll show you x and y in terms of a and b.

OpenStudy (anonymous):

@Hero @SolomonZelman Please help?

OpenStudy (kainui):

If you're just going to ignore me... why do you think they'll help you any more than me who's here trying to help?

OpenStudy (anonymous):

Yeah but I don't know what a matrix equation is, I'm sorry for ignoring you.

OpenStudy (anonymous):

Sorry, I should have told you that @Kainui

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Hey there! Are you able to help me? lol.

OpenStudy (anonymous):

multiply the second equation by 2 to get this -2x + 10y = 12b then add the 2 equations together like this 2x + 3y =a -2x + 10y = 12b ============ to get this 0 + 13y = a + 12b from here you find that y = (a+12b)/13 now substitute this y into any of the 2 original equation to find your x good luck

OpenStudy (anonymous):

showed you above another way to find the answer if you dont know yet how to work with matrices, perhaps your class hasnt covered it yet...